test. Please keep your answers in two decimal points forl P 1. Joes Camera to Yols Camera shop has a favorite model that has annual sales of 400. The cost to place replenish inventory is $40 per order, and annual holding cost is $20 per unit per year store is open 300 days per year. Assume the What is the optimal order size? a. b. What is the optimal number of orders per year? c. What is the optimal number of days between orders? d. What is the annual cost of ordering and holding inventory?
Answer #1

Annual demand (D) = 400 units

Ordering cost (S) = $40 per order

Holding cost (H) = $20

Number of days per year = 300 days

a) Optimal order size (Q) = sqrt of (2DS / H)

= sqrt of [(2 x 400 x 40) /20]

= sqrt of 1600

= 40 units

b) Number of orders per year = D/Q = 400/40 = 10 orders

C) Time between orders = (Q/D)Number of days per year

= (40/400)300

= (0.1x300)

= 30 days

d) Annual ordering cost = (D/Q) S = (400/40)40 = $400

Annual holding cost = (Q/2) H = (40/2)20 = $400

Annual cost of ordering and holding inventory = ordering cost + Holding cost = $400+$400 = $800